# Algorithm for finding local integrals of motion in quantum lattice models in the thermodynamic limit

### v1

@@ Abstract
Local integrals of motion (LIOMs) play a key role in understanding the stationary states of closed macroscopic systems. They were found for selected integrable systems via complex analytical calculations. The existence of LIOMs and their structure can also be studied via numerical methods, which, however, involve exact diagonalization of Hamiltonians, posing a bottleneck for such studies. We show that finding LIOMs in translationally invariant lattice models or unitary quantum circuits can be reduced to a problem for which one may numerically find an exact solution also in the thermodynamic limit. We develop and implement a simple algorithm and demonstrate the efficiency of this method by calculating LIOMs and the Mazur bounds for infinite integrable spin chains and unitary circuits. Finally, we demonstrate that this approach correctly identifies approximate LIOMs in nearly integrable spin ladders.
@@
Introduction – Closed macroscopic systems during time-evolution reach a stationary state that is determined by local conservation laws or, equivalently, by local integrals of motion (LIOMs). Generic systems have only a few conserved quantities (like Hamiltonian, particle number, magnetization) and evolve to a thermal Gibbs state [1, 2, 3]. Interacting integrable systems have extensive number of LIOMs and the relevant stationary states have the form that is consistent with the generalized Gibbs ensemble [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. Identification of LIOMs is also the starting point for construction of the hydrodynamics or the generalized hydrodynamics [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31] in case of integrable models. Moreover, anomalous transport (or its anticipated absence) in various quantum systems has recently been linked to the presence of LIOMs. Most prominent examples concern strongly disordered systems [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43], tilted chains with emergent conservation of the dipole moment [44] and systems exhibiting the Hilbert space fragmentation [45, 46, 44].
@@
While the importance of LIOMs for the long-time dynamics appears unquestionable, in many cases their construction remains a challenging problem. Even for the most studied integrable interacting model, i.e. the XXZ chain, the complete set of LIOMs (that includes also quasilocal LIOMs) has been established analytically only recently [47, 48, 49, 50, 13, 51, 52]. In many cases, LIOMs are identified within complex calculations that are specific for the given studied model [53, 54, 55, 56, 57, 58, 59, 60]. Although there are model-independent approaches which allow to find LIOMs in integrable models [61, 62] or approximate LIOMs in nearly integrable systems [63], they are based on exact diagonalization of Hamiltonians. The latter poses serious limitations on the studied models and raises questions concerning the reliability of the finite-size scaling.
@@
In this Letter, we show that finding LIOMs in translationally-invariant lattice models belongs to the same class of problems as the high-temperature expansion [64, 65, 66]. Namely, one can numerically find exact LIOMs and the corresponding Mazur bounds [67, 68] in the thermodynamic limit without diagonalizing the Hamiltonian. This holds true for quantum systems where the dynamics is set either by a Hamiltonian or by a unitary transformation, as is the case in unitary circuits. The geometry of the system is not important, and the only essential assumption concerns the dynamics that needs to be set by local transformations.
@@
Method – We consider a system containing $L$ sites, which are enumerated by the index $l=1,\ldots,L$, and study extensive Hermitian operators, $\hat{O}^{s}=\sum_{l}\hat{O}^{s}_{l}$, where the upper index, $s$, distinguishes between various operators. The densities, $\hat{O}^{s}_{l}$, are local and act on up to $M$ sites close to $l$. We introduce the Hilbert-Schmidt (HS) product $\langle\hat{O}^{s}\hat{O}^{p}\rangle={\rm Tr}(\hat{O}^{s\dagger}\hat{O}^{p})/(ZL)$ and the HS norm of operators, $||\hat{O}^{s}||^{2}=\langle\hat{O}^{s}\hat{O}^{s}\rangle$, where $Z$ is the dimension of the Hilbert space. We recall that the HS products of local operators, $\langle\hat{O}^{s}_{l}\hat{O}^{s^{\prime}}_{l^{\prime}}\rangle$, can be exactly calculated for infinite lattice systems using the technique employed in the high-temperature expansion.
@@
Our main assumption is that the dynamics is set either by a local Hamiltonian, $\hat{H}$, or by a repeated sequence of local unitary transformations $\hat{U}$. In the former case, we take $\hat{H}=\sum_{l}\hat{H}_{l}$, where $\hat{H}_{l}$ acts on $M_{H}$ sites. We look for local operators, $\hat{Q}$, which commute with the Hamiltonian, $[\hat{Q},\hat{H}]=0$ or, equivalently [69], we require vanishing of the norm of a Hermitian operator $\hat{Q}^{\prime}=i[\hat{Q},\hat{H}]$, i.e., $||\hat{Q}^{\prime}||^{2}=\langle\hat{Q}^{\prime}\hat{Q}^{\prime}\rangle=0$. We note that $\hat{Q}^{\prime}$ is still local, $\hat{Q}^{\prime}=\sum_{l}\hat{Q}^{\prime}_{l}$, where $\hat{Q}^{\prime}_{l}$ acts on up to $M+M_{H}-1$ sites. For this reason, $\langle\hat{Q}^{\prime}\hat{Q}^{\prime}\rangle$ can also be calculated exactly in the thermodynamic limit. In case of circuits built out of repeated sequence of local unitary transformations, the same reasoning can be applied for $\hat{Q}^{\prime}=\hat{U}^{\dagger}\hat{Q}\hat{U}-\hat{Q}$. I.e., $\hat{Q}^{\prime}$ is a local operator and vanishing norm of $\hat{Q}^{\prime}$ implies that $\hat{Q}$ is conserved. Both cases can be comprehended within a unified formulation. Namely, we introduce a function $f(\hat{O}^{s})$ which has the following properties: (i) $f$ is linear, (ii) vanishing of $||f(\hat{O}^{s})||$ is equivalent to conservation of $\hat{O}^{s}$; (iii) $f(\hat{O}^{s})$ is a local and Hermitian operator so that one may exactly calculate $\langle f(\hat{O}^{s})f(\hat{O}^{p})\rangle$. If the dynamics is set by a Hamiltonian then one may choose $f(\hat{O}^{s})=i[\hat{O}^{s},\hat{H}]$, whereas $f(\hat{O}^{s})=\hat{U}^{\dagger}\hat{O}^{s}\hat{U}-\hat{O}^{s}$ can be used for unitary circuits. The specific choice of $f$ is irrelevant as long as it is consistent with conditions (i)-(iii).
@@
As a next step, we consider a set containing $N_{O}$ orthonormal operators, $\langle\hat{O}^{s}\hat{O}^{p}\rangle=\delta_{sp}$. We aim to find all LIOMs which can be expressed as linear combinations of these operators. To this end, we build the matrix containing HS products $F_{sp}=\langle f(\hat{O}^{s})f(\hat{O}^{p})\rangle$ and solve the eigenproblem,
@@
where $V$ is an orthogonal matrix and $F$ is a matrix built in the basis (of size $N_{O}$) containing all $\hat{O}^{s}$. Since $f$ is a linear function, one may introduce rotated operators
@@
from which one finds the norm $||f(\hat{A}^{\alpha})||=\sqrt{\lambda_{\alpha}}$. The rotated operators are local and orthogonal, $\langle\hat{A}^{\alpha}\hat{A}^{\beta}\rangle=\delta_{\alpha\beta}$. Consequently, all $\hat{A}^{\alpha}$ corresponding to the eigenvalue $\lambda_{\alpha}=0$ represent orthogonal LIOMs, identified here via the norm $||f(\hat{A}^{\alpha})||=0$. To distinguish these LIOMs from other operators introduced in Eq. (2), we denote LIOMs as $\hat{Q}^{\alpha}$ where $\alpha=1,\dots,N_{L}$ and $N_{L}$ is the number of LIOMs given by the degeneracy of the eigenvalue $\lambda_{\alpha}=0$ in Eq. (1). We keep the notation $\hat{A}^{\alpha}$, for the remaining operators, $\alpha=N_{L}+1,...,N_{O}$ with $\lambda_{\alpha}>0$.
As a final step, we show that the solution of Eq. (1) identifies all LIOMs that can be represented via linear combinations of $\hat{O}^{s}$. To this end, we consider an operator $\hat{B}=\sum_{s}b_{s}\hat{O}^{s}$, $||\hat{B}||>0$, that is orthogonal to all $\hat{Q}^{\alpha}$ obtained from Eq.(1). We show that $||f(\hat{B})||>0$ implying that $\hat{B}$ is not conserved. Using the transformation from Eq. (2), such operator can be expressed in the rotated basis
$\hat{B}$ $=$ $\sum_{\alpha=1}^{N_{L}}\left(\sum_{s}b_{s}V_{\alpha s}\right)\hat{Q}^{\alpha}+\sum_{\alpha=N_{L}+1}^{N_{O}}\left(\sum_{s}b_{s}V_{\alpha s}\right)\hat{A}^{\alpha}$ (4)
$=$ $\sum_{\alpha=1}^{N_{L}}\tilde{b}_{\alpha}\hat{Q}^{\alpha}+\sum_{\alpha=N_{L}+1}^{N_{O}}\tilde{b}_{\alpha}\hat{A}^{\alpha}.$
The first term in Eq. (4) vanishes since we assume that $\hat{B}$ is orthogonal to all $\hat{Q}^{\alpha}$. Then, one finds the squared norm
$||f(\hat{B})||^{2}$ $=$ $\sum_{\alpha,\beta=N_{L}+1}^{N_{O}}\tilde{b}_{\alpha}\tilde{b}_{\beta}\langle f(\hat{A}^{\alpha})f(\hat{A}^{\beta})\rangle$ (5)
$=$ $\sum_{\alpha=N_{L}+1}^{N_{O}}\tilde{b}^{2}_{\alpha}\lambda_{\alpha}>0,$
where we have used Eq. (3). The inequality in Eq. (5) follows from that all $\lambda_{\alpha}$ are positive for $\alpha>N_{L}$ and some $\tilde{b}_{\alpha}$ are nonzero (otherwise $||\hat{B}||=0$, contradicting one of the assumptions).
Solving the eigenproblem in Eq. (1) gives not only the LIOMs but also the Mazur bounds for all operators $\hat{O}^{s}$ used in the construction of the $F$ matrix. Using Eq.(2) one finds that the infinite-time ($t$) correlation functions are bounded by the matrix $V$,
@@
The basis of operators in spin-1/2 systems – The first step is to choose the set of operators, $\hat{O}_{s}$, for which we then construct the matrix in Eq. (1). Any specific knowledge concerning LIOMs may help reduce the set of operators that are used for its construction. Without such knowledge, one may implement a brute-force approach and systematically include all operators with growing support. In such case, one may use at each site, $l$, the set containing four operators: the identity matrix, $\hat{\sigma}^{0}_{l}$, and the Pauli matrices $\hat{\sigma}^{x}_{l}$, $\hat{\sigma}^{y}_{l}$, $\hat{\sigma}^{z}_{l}$. These operators are orthogonal, $\langle\hat{\sigma}^{s}_{l}\hat{\sigma}^{s^{\prime}}_{l}\rangle=\delta_{s,s^{\prime}}/L$, so they may be used for constructing extensive operators, $\hat{O}^{s}=\sum_{l}\hat{O}^{s}_{l}$, supported on $m$ sites
$\hat{O}^{s}_{l}=\hat{O}^{\left(s_{1},...,s_{m}\right)}_{l}=\hat{\sigma}^{s_{1}}_{l+1}\hat{\sigma}^{s_{2}}_{l+2}...\hat{\sigma}^{s_{m}}_{l+m}.$ (7)
@@
We assume that the first, $\hat{\sigma}^{s_{1}}_{l+1}$, and the last, $\hat{\sigma}^{s_{m}}_{l+m}$, spin operators are other than $\hat{\sigma}^{0}$. All operators with maximal support $M$ (i.e., generated for all $m\leq M$) are orthogonal $\langle\hat{O}^{s}_{l}\hat{O}^{s^{\prime}}_{l^{\prime}}\rangle=\delta_{l,l^{\prime}}\delta_{s,s^{\prime}}/L$. For fixed $M$, one has in total $N_{O}=3\times 4^{M-1}$ extensive operators which are orthonormal $\langle\hat{O}^{s}\hat{O}^{s^{\prime}}\rangle=\delta_{s,s^{\prime}}$. We note that all translationally invariant operators supported on up to $M$ sites can be expressed as linear combinations of $\hat{O}^{s}$ thus one may easily calculate the HS product of such operators. Finally, we recall that product of two $\hat{\sigma}^{s}_{l}$-operators is proportional to other such operator, $\hat{\sigma}^{s}_{l}\hat{\sigma}^{s^{\prime}}_{l}=\alpha_{s^{\prime\prime}}\hat{\sigma}^{s^{\prime\prime}}_{l}$, where $\alpha_{s^{\prime\prime}}\in\{1,\pm i\}$. This enables a straightforward calculation of $f(\hat{O}^{s})$, which we implement using the software package from Ref. [70].

### v2

# Finding local integrals of motion in quantum lattice models in the thermodynamic limit

@@ Abstract
Local integrals of motion (LIOMs) play a key role in understanding the long-time properties of closed macroscopic systems. They were found for selected integrable systems via complex analytical calculations. The existence of LIOMs and their structure can also be studied via numerical methods, which, however, involve exact diagonalization of Hamiltonians, posing a bottleneck for such studies. We show that finding LIOMs in translationally invariant lattice models or unitary quantum circuits can be reduced to a problem for which one may numerically find an exact solution in the thermodynamic limit. We develop a simple algorithm and demonstrate its efficiency by calculating LIOMs and bounds on correlations (the Mazur bounds) for infinite integrable spin chains and unitary circuits. Finally, we demonstrate that this approach identifies slow modes in nearly integrable spin models and estimates their relaxation times.
I Introduction
@@
Closed macroscopic systems during time-evolution reach a stationary state that is determined by local conservation laws or, equivalently, by local integrals of motion (LIOMs). Generic systems have only a few conserved quantities (like Hamiltonian, particle number, magnetization) and evolve to a thermal Gibbs state [1, 2, 3]. Interacting integrable systems have extensive number of LIOMs and the relevant stationary states have the form that is consistent with the generalized Gibbs ensemble [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. Identification of LIOMs is also the starting point for construction of the hydrodynamics or the generalized hydrodynamics [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31] in case of integrable models. Moreover, anomalous transport (or its anticipated absence) in various quantum systems has recently been linked to the presence of LIOMs. Most prominent examples concern strongly disordered systems [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43], tilted chains with emergent conservation of the dipole moment [44] and systems exhibiting the Hilbert space fragmentation [45, 46, 44].
@@
While the importance of LIOMs for the long-time dynamics appears unquestionable, in many cases their construction remains a challenging problem. Even for the most studied integrable interacting model, i.e., the XXZ chain, the complete set of LIOMs (that includes also quasilocal LIOMs) has been established analytically only recently [47, 48, 49, 50, 13, 51, 52]. In many cases, LIOMs are identified through complex calculations that are specific to the investigated model [53, 54, 55, 56, 57, 58, 59, 60]. Although there are model-independent approaches which allow to find LIOMs in integrable models [61, 62] or approximate LIOMs in nearly integrable systems [63], they are based on exact diagonalization of Hamiltonians. The latter poses serious limitations on the studied models and raises questions concerning the reliability of the finite-size scaling.
@@
In this work, we show that finding LIOMs in translationally-invariant lattice models belongs to the same class of problems as the high-temperature expansion [64, 65, 66]. Namely, one can numerically find exact LIOMs and the corresponding Mazur bounds [67, 68] in the thermodynamic limit without diagonalizing the Hamiltonian. This holds true for quantum systems where the dynamics is set either by a Hamiltonian or by a unitary transformation, as is the case in unitary circuits. The geometry of the system is not important, and the only essential assumption concerns the dynamics that needs to be set by local transformations.
II General approach
II.1 Algorithm
@@
We consider a system containing $L$ sites, which are enumerated by the index $l=1,\ldots,L$, and study extensive Hermitian operators, $\hat{O}^{s}=\sum_{l}\hat{O}^{s}_{l}$, where the upper index, $s$, distinguishes between various operators. The densities, $\hat{O}^{s}_{l}$, are local and act on up to $M$ sites adjacent to $l$. We introduce the Hilbert-Schmidt (HS) product $\langle\hat{O}^{s}\hat{O}^{p}\rangle={\rm Tr}(\hat{O}^{s\dagger}\hat{O}^{p})/(ZL)$ and the HS norm of operators, $\lVert\hat{O}^{s}\lVert^{2}=\langle\hat{O}^{s}\hat{O}^{s}\rangle$, where $Z$ is the dimension of the Hilbert space. We recall that the HS product of the local densities, $\langle\hat{O}^{s}_{l}\hat{O}^{s^{\prime}}_{l^{\prime}}\rangle$, can be exactly calculated for infinite lattice systems using just the properties of the operator algebra, similarly to the technique employed in the high-temperature expansion.
@@
Our main assumption is that the dynamics is set either by a local Hamiltonian, $\hat{H}$, or by a repeated sequence of local unitary transformations $\hat{U}$. In the former case, we take $\hat{H}=\sum_{l}\hat{H}_{l}$, where $\hat{H}_{l}$ acts on $M_{H}$ sites. We look for local operators, $\hat{Q}$, which commute with the Hamiltonian, $[\hat{Q},\hat{H}]=0$ or, equivalently, we require vanishing of the norm of a Hermitian operator $\hat{Q}^{\prime}=i[\hat{Q},\hat{H}]$, i.e., $\lVert\hat{Q}^{\prime}\lVert^{2}=\langle\hat{Q}^{\prime}\hat{Q}^{\prime}\rangle=0$. We note that $\hat{Q}^{\prime}$ is still local, $\hat{Q}^{\prime}=\sum_{l}\hat{Q}^{\prime}_{l}$, where $\hat{Q}^{\prime}_{l}$ acts on up to $M+M_{H}-1$ sites. For this reason, $\langle\hat{Q}^{\prime}\hat{Q}^{\prime}\rangle$ can also be calculated exactly in the thermodynamic limit. In case of circuits built out of repeated sequence of local unitary transformations, the same reasoning can be applied for $\hat{Q}^{\prime}=\hat{U}^{\dagger}\hat{Q}\hat{U}-\hat{Q}$. I.e., $\hat{Q}^{\prime}$ is a local operator and vanishing norm of $\hat{Q}^{\prime}$ implies that $\hat{Q}$ is conserved. Both cases can be comprehended within a unified formulation. Namely, we introduce a function $f(\hat{O}^{s})$ which has the following properties: (i) $f$ is linear, (ii) vanishing of $\lVert f(\hat{O}^{s})\lVert$ is equivalent to conservation of $\hat{O}^{s}$; (iii) $f(\hat{O}^{s})$ is a local and Hermitian operator so that one can exactly calculate $\langle f(\hat{O}^{s})f(\hat{O}^{p})\rangle$. If the dynamics is set by a Hamiltonian then one may choose $f(\hat{O}^{s})=i[\hat{O}^{s},\hat{H}]$, whereas $f(\hat{O}^{s})=\hat{U}^{\dagger}\hat{O}^{s}\hat{U}-\hat{O}^{s}$ can be used for unitary circuits. The specific choice of $f$ is irrelevant and can be adapted to the desired setup, as long as it is consistent with conditions (i)-(iii).
@@
As a next step, we consider a set containing $N_{O}$ orthonormal operators, $\langle\hat{O}^{s}\hat{O}^{p}\rangle=\delta_{sp}$. We aim to find all LIOMs which can be expressed as linear combinations of these operators. To this end, we solve the eigenproblem
@@
where $V$ is an orthogonal matrix and $F$ is a matrix of HS products, $F_{sp}=\langle f(\hat{O}^{s})f(\hat{O}^{p})\rangle$, constructed in the basis (of size $N_{O}$) containing all $\hat{O}^{s}$. Since $f$ is a linear function, one may introduce rotated operators
@@
from which one finds the norm $\lVert f(\hat{A}^{\alpha})\lVert=\sqrt{\lambda_{\alpha}}$. The rotated operators are local and orthogonal, $\langle\hat{A}^{\alpha}\hat{A}^{\beta}\rangle=\delta_{\alpha\beta}$. Consequently, all $\hat{A}^{\alpha}$ corresponding to the eigenvalue $\lambda_{\alpha}=0$ represent orthogonal LIOMs, identified here via the norm $\lVert f(\hat{A}^{\alpha})\lVert=0$. To distinguish these LIOMs from other operators introduced in (2), we denote LIOMs as $\hat{Q}^{\alpha}$ where $\alpha=1,\dots,N_{L}$ and $N_{L}$ is the number of LIOMs given by the degeneracy of the eigenvalue $\lambda_{\alpha}=0$ in (1). We keep the notation $\hat{A}^{\alpha}$, for the remaining operators, $\alpha=N_{L}+1,...,N_{O}$ with $\lambda_{\alpha}>0$.
As a final step, we show that the solution of (1) identifies all LIOMs that can be represented via linear combinations of $\hat{O}^{s}$. To this end, we consider an operator $\hat{C}=\sum_{s}c_{s}\hat{O}^{s}$, $\lVert\hat{C}\lVert>0$, that is orthogonal to all $\hat{Q}^{\alpha}$ obtained from (1). We show $\lVert f(\hat{C})\lVert>0$, implying that $\hat{C}$ is not conserved. Using the inverse of the transformation from (2), such operator can be expressed in the rotated basis
$\hat{C}$ $=$ $\sum_{\alpha=1}^{N_{L}}\left(\sum_{s}c_{s}V_{\alpha s}\right)\hat{Q}^{\alpha}+\sum_{\alpha=N_{L}+1}^{N_{O}}\left(\sum_{s}c_{s}V_{\alpha s}\right)\hat{A}^{\alpha}$ (4)
$=$ $\sum_{\alpha=1}^{N_{L}}\tilde{c}_{\alpha}\hat{Q}^{\alpha}+\sum_{\alpha=N_{L}+1}^{N_{O}}\tilde{c}_{\alpha}\hat{A}^{\alpha}.$
The first term in (4) vanishes since we assume that $\hat{C}$ is orthogonal to all $\hat{Q}^{\alpha}$. Then, one finds the squared norm
$\lVert f(\hat{C})\lVert^{2}$ $=$ $\sum_{\alpha,\beta=N_{L}+1}^{N_{O}}\tilde{c}_{\alpha}\tilde{c}_{\beta}\langle f(\hat{A}^{\alpha})f(\hat{A}^{\beta})\rangle$ (5)
$=$ $\sum_{\alpha=N_{L}+1}^{N_{O}}\tilde{c}^{2}_{\alpha}\lambda_{\alpha}>0,$
where we have used (3). The inequality in (5) follows from that all $\lambda_{\alpha}$ are positive for $\alpha>N_{L}$ and some $\tilde{c}_{\alpha}$ are nonzero (otherwise $\lVert\hat{C}\lVert=0$, contradicting one of the assumptions).
Solving the eigenproblem in (1) gives not only the LIOMs but also the Mazur bounds for all operators $\hat{O}^{s}$ used in the construction of the $F$ matrix. Using Eq.(2) one finds that the infinite-time ($t$) correlation functions are bounded by the matrix $V$,
@@
For systems where the dynamics is set by the Hamiltonian, our approach is consistent with the idea from Ref. [69, 70, 71] based on minimizing the norm of commutators $[\hat{H},\hat{A}]$. For quantum circuits, it is similar (but not equivalent) to the approach from Refs. [72, 73, 74, 75] based on Ruelle-Pollicott resonances obtained from diagonalization of a non-Hermitian matrix $\langle f(\hat{O}^{s})\hat{O}^{p}\rangle$.
II.2 The basis of operators
@@
The first step in our algorithm is to choose the set of operators, $\hat{O}_{s}$, for which we then construct the matrix in (1). Any specific knowledge concerning LIOMs may help reduce the set of operators that are used for its construction. Without such knowledge, one may implement a brute-force approach and systematically include all operators with growing support. In such case, one may use at each site, $l$, the set containing four operators: the identity matrix, $\hat{\sigma}^{0}_{l}$, and the Pauli matrices $\hat{\sigma}^{x}_{l}$, $\hat{\sigma}^{y}_{l}$, $\hat{\sigma}^{z}_{l}$. These operators are orthogonal, $\langle\hat{\sigma}^{s}_{l}\hat{\sigma}^{s^{\prime}}_{l}\rangle=\delta_{s,s^{\prime}}/L$, so they may be used for constructing extensive operators, $\hat{O}^{s}=\sum_{l}\hat{O}^{s}_{l}$, supported on $m$ sites
$\hat{O}^{s}_{l}=\hat{O}^{\left(s_{1},...,s_{m}\right)}_{l}=\hat{\sigma}^{s_{1}}_{l}\hat{\sigma}^{s_{2}}_{l+1}...\hat{\sigma}^{s_{m}}_{l+m-1}.$ (7)
@@
In the following, we assume that the first, $\hat{\sigma}^{s_{1}}_{l}$, and the last, $\hat{\sigma}^{s_{m}}_{l+m-1}$, spin operators are other than $\hat{\sigma}^{0}$. All operators with maximal support $M$ (i.e., generated for all $m\leq M$) are orthogonal $\langle\hat{O}^{s}_{l}\hat{O}^{s^{\prime}}_{l^{\prime}}\rangle=\delta_{l,l^{\prime}}\delta_{s,s^{\prime}}/L$. For fixed $M$, one has in total $N_{O}=3\times 4^{M-1}$ extensive operators which are orthonormal $\langle\hat{O}^{s}\hat{O}^{s^{\prime}}\rangle=\delta_{s,s^{\prime}}$. We note that all translationally invariant operators supported on up to $M$ sites can be expressed as linear combinations of $\hat{O}^{s}$, thus one may easily calculate the HS product of such operators. Finally, we recall that product of two $\hat{\sigma}^{s}_{l}$-operators is proportional to other such operator, $\hat{\sigma}^{s}_{l}\hat{\sigma}^{s^{\prime}}_{l}=\alpha_{s^{\prime\prime}}\hat{\sigma}^{s^{\prime\prime}}_{l}$, where $\alpha_{s^{\prime\prime}}\in\{1,\pm i\}$. This enables a straightforward calculation of $f(\hat{O}^{s})$, which we implement using the software package from Ref. [76].
III Results
III.1 LIOMs in the XXZ chain